# Attributed-graphs kernel implementation using local detuning of neutral-atoms Rydberg Hamiltonian

### v1

@@ III Graph Kernel methods for machine learning
In this section we introduce all the kernels used in this work: first, the classical kernels used for benchmarking; then, the quantum feature kernels, QEK (III.1) and GDQC (III.2). Lastly, in Subsec. III.3 we introduce the pooling operations that are employed.
@@
Kernel-based machine learning methods rely on a rigorous mathematical framework grounded in the theory of Reproducing Kernel Hilbert Spaces (RKHS). The foundation is laid by the Moore-Aronszajn Theorem, which guarantees that every positive definite kernel defines a unique RKHS where the kernel acts as the reproducing kernel [28] . Within this space, the Representer Theorem shows that the solution to many regularized risk minimization problems can be written as a finite linear combination of kernel functions evaluated at the training points, thus reducing infinite-dimensional optimization to a tractable finite-dimensional one [29]. Furthermore, Mercer’s Theorem [30] offers a spectral decomposition of continuous, symmetric, positive-definite kernels, interpreting them as inner products in high-dimensional feature spaces [31]. These results together enable the development of powerful nonlinear algorithms, such as Support Vector Machines (SVMs) [32], with solid theoretical guarantees and practical scalability, for various machine learning tasks such as classification and regression. In practical terms these kernel methods allow SVMs to classify non-linearly separable data by implicitly mapping input features into a higher-dimensional space where a linear separator can be found, without the need of an explicit computation of the mapping [31].
Graph-based data are known to be a particular kind of data due to the non-Euclidean nature of their associated distance measure. This limits the scope of the types of kernels that can be used on graphs by putting additional constraint on their properties (see Sec. III.2 for more details). Therefore, researchers came up with graph-tailored kernels that are typically designed to overcome these limitations by capturing structural similarities. This is achieved, for example, by means of walks, subtrees, or subgraph patterns. By embedding graphs into Hilbert spaces while preserving their combinatorial properties, they provide a principled and flexible approach to graph machine learning tasks. From the vast literature of classical graph kernels, we employ the following ones for the benchmarking purposes:
Random walk kernels.
These measure similarity between graphs by counting matching random walks in each graph, capturing detailed structural patterns. Introduced in Ref. [33], they offer expressiveness but can be computationally intensive and prone to tottering effects.
k-core kernels.
These kernels measure similarity between graphs by comparing the distributions of their $k$-cores, which are defined as the maximal number of subgraphs where every node has a degree at least equal to $k$. The approach was proposed in Ref. [34] as part of structured representations for graphs.
Shortest path kernel.
Introduced in Ref. [35] this kernel is an efficient alternative to random walks, capturing both local and global structures by evaluating pairs of nodes with matching shortest path lengths.
Weisfeiler-Leman kernels.
Based on the Weisfeiler and Leman (WL) graph isomorphism test, these kernels iteratively update node labels based on neighborhood structures [6].
III.1 Quantum evolution kernel (QEK)
@@
First introduced by Henry et al. [13], the quantum evolution kernel (QEK) algorithm is a quantum-feature kernel implementation [see Eq. (1)] that draws from probability distributions generated from measured observables. Specifically, within the proof-of-principle implementation of Ref. [14], the algorithm draws from the global observable vector $\bm{P}^{\mathcal{G}}=\left(P_{0}^{\mathcal{G}}(t),P_{1}^{\mathcal{G}}(t),P_{2}^{\mathcal{G}}(t),\dots\right)^{\intercal}$ with $P_{k}^{\mathcal{G}}(t)$ being the excitation probability for a fixed number $k$ of excitations at a given time $t$, see Eq. (6). The superscript $\mathcal{G}$ indicates that the time evolution is subjected to a graph-embedded Hamiltonian, cf. Sec. II. To discriminate between two graphs, $\mathcal{G}_{\mu}$ and $\mathcal{G}_{\mu^{\prime}}$, the QEK algorithm implements the following kernel function
$\kappa_{\mathrm{QEK}}(\mathcal{G}_{\mu},\mathcal{G}_{\mu^{\prime}})=\exp\left[-\mu_{0}\,\mathrm{JS}\left(\bm{P}^{\mathcal{G_{\mu}}},\bm{P}^{\mathcal{G_{\mu^{\prime}}}}\right)\right].$ (12)
@@
Here, the Jensen–Shannon divergence $\mathrm{JS}\left(\bm{P},\bm{P}^{\prime}\right)$ measures the similarity between the two distributions. By construction, the kernel function (12) ensures positive definiteness. Throughout this work we fix the model parameter $\mu_{0}=2$.
@@
Within its proof-of-principle implementation [14], the QEK kernel (12) has been applied to non-attributed graphs. Correspondingly, the time evolution has been subjected to a global-detuning Hamiltonian $\hat{\mathcal{H}}_{\mathrm{glob}}^{\mathcal{G}}$ (9). When coupled with a support vector machine (SVM), the QEK algorithm has matched the performance of leading state-of-the-art classical graph kernels on several benchmark datasets of small graphs [14]. Thereby, QEK‘s capabilities have been demonstrated both, numerically via emulations as well as on real quantum hardware, on a 32-qubit neutral-atom processor.
@@
In our work, we expand on the previous QEK implementation [14] twofold. First, and most importantly, we want it to be able to deal with attributed graphs. This is accomplished by changing the Hamiltonian that governs the time evolution from $\hat{\mathcal{H}}_{\mathrm{glob}}^{\mathcal{G}}$ (9) to the local-detuning Hamiltonian $\hat{\mathcal{H}}_{\mathrm{loc}}^{\mathcal{G}}$ (10). As described in Subsec. II.2, the latter Hamiltonian embeds the atomic masses (the node features) into the local detuning $\delta_{i}$. As a result, we obtain an observable vector $\bm{P}^{\mathcal{G}}$ which contains graph information drawn from both, the edges and the node features. Similar to before, the observable vectors are inserted into the kernel function (12). To distinguish the two cases, evolution under $\hat{\mathcal{H}}_{\mathrm{glob}}^{\mathcal{G}}$ and $\hat{\mathcal{H}}_{\mathrm{loc}}^{\mathcal{G}}$, we use the notation $\kappa_{\mathrm{QEK}}^{\mathrm{glob}}$ and $\kappa_{\mathrm{QEK}}^{\mathrm{loc}}$, respectively. Mathematically, switching from $\kappa_{\mathrm{QEK}}^{\mathrm{glob}}$ to $\kappa_{\mathrm{QEK}}^{\mathrm{loc}}$ has profound consequences which can be summarized by proposition 1 which is proven in App. A.
@@
The quantum evolution kernel $\kappa_{\mathrm{QEK}}^{\mathrm{loc}}$ is genuinely more expressive than $\kappa_{\mathrm{QEK}}^{\mathrm{glob}}$ for dinstinguishing attributed graphs.
@@
The second extension to the QEK algorithm that we undertake lies in pooling of different time steps. The above implementation draws the excitation probability from a single time step $t$. In Subsec. III.3 we explain how multiple time steps can be united in order to combine information sampled from different regimes of the time evolution.
@@
The excitation probability $P_{k}^{\mathcal{G}}(t)$ (6) used in the QEK algorithm is a natural observable candidate to capture graph information as it is permutation invariant. Moreover, it is a global observable rendering it a useful tool to study global graph properties. However, if it comes to local graph properties the excitation probability $P_{k}^{\mathcal{G}}(t)$ may not be the most convenient choice. An attempt to generalize the algorithm from the global observables $P_{k}^{\mathcal{G}}(t)$ to observables that capture distributions of local subcomponents of a graph (e.g., nodes or edges)—while maintaining permutation invariance—requires the fixing of a canonical ordering for each graph, a costly task known as the graph canonization problem [36]. This potential shortcoming of the QEK algorithm (12) has motivated the design of the generalized-distance quantum-correlation (GDQC) kernel, see Subsec. III.2.
III.2 Generalized-distance quantum-correlation (GDQC) kernel
@@
In this part we introduce the new quantum-feature kernel which is used in this work, the generalized-distance quantum-correlation (GDQC) kernel. In contrast to the QEK algorithm which designed to excel at capturing global graph information, the new kernel draws from local observables making it more apt at predicting local graph properties.
The GDQC kernel is inspired by the classical generalized-distance Weisfeiler-and-Leman (GD-WL) kernel [6]. As such the goal is to combine the graph distance, i.e. distance $d_{\mathcal{G}}(i,j)$ between nodes $i$ and $j$ and $D^{\mathcal{G}}$ being the distance matrix where $D^{\mathcal{G}}_{i,j}=d_{\mathcal{G}}(i,j)$, with another piece of graph information. Within the GDQC framework, the second piece of information is a local observable. In the present work we use the correlation matrix $C^{{}^{\mathcal{G}}}(t)$ where $C^{\mathcal{G}}_{i,j}\left(t\right)=\left\langle\psi_{\mathcal{G}}\left(t\right)\big{|}\hat{n}_{i}\hat{n}_{j}\big{|}\psi_{\mathcal{G}}\left(t\right)\right\rangle$, cf. Eq. (6). However, it should be noted that the GDQC framework can be easily generalized to work with other local observables, including higher-order ones.
@@
Algorithm 1 Computation of the GDQC feature vector $\bm{\chi}\!\left[C^{\mathcal{G}}\!\left(t\right),D^{\mathcal{G}}\right]$ in Eq. (13).
The proposed GDQC kernel function [cf. Eq. (1)] between two graphs, $\mathcal{G}_{\mu}$ and $\mathcal{G}_{\mu^{\prime}}$, is defined as scalar product,
@@
It involves a graph-dependent GQDC feature vector $\bm{\chi}\left[C^{\mathcal{G}}\left(t\right),D^{\mathcal{G}}\right]$ whose construction is the core element of the GDQC framework. Its construction is mathematically defined via algorithm 1, and it is illustrated in Fig. 2. It relies on binning procedures where separately the correlation matrix $C^{\mathcal{G}}_{i,j}\left(t\right)\in\left[0,1\right]$ and the distance matrix $D^{\mathcal{G}}_{i,j}$ are binned. The former is binned into $N_{\mathrm{bins}}^{C}\geq 1$ intervals of size $\varDelta_{C}=1/N_{\mathrm{bins}}^{C}$ (bins associated with the upper axis in Fig.  2, while the latter is binned into $N_{\mathrm{bins}}^{D}\geq 1$ intervals of size $\varDelta_{D}=d_{\mathrm{max}}/N_{\mathrm{bins}}^{D}$ where $d_{\mathrm{max}}$ denotes the maximum distance across the entire dataset. In our case we work with graph distance (or the shortest-path distance), i.e. the distance measure $d_{\mathcal{G}}(i,j)\in\mathbb{N}$ is integer, which simplifies the process of distance binning. Here, the most natural choice is to use $N_{\mathrm{bins}}^{D}=d_{\mathrm{max}}+1$ bins, i.e. one bin for each value of graph distance (including distance 0) as illustrated on the lower axis in Fig. 2. Concerning the correlation binning we study the impact of the binning size $\varDelta_{C}=1/N_{\mathrm{bins}}^{C}$ in Subsec. IV.1. The results presented in Sec. IV are derived for $N_{\mathrm{bins}}^{C}=10$. In combination the two binning procedures create the $(N_{\mathrm{bins}}^{C}N_{\mathrm{bins}}^{D})$-dimensional GDQC feature vector $\bm{\chi}$ where each element is associated with a specific combination of correlation and distance bins. The value of each element is given as the count of correlation matrix elements $C^{\mathcal{G}}_{i,j}\left(t\right)$ falling into the right value range, and whose involved nodes fall into the respective distance bin. In Fig. 2 we present a simple example of this binning procedure. In our study, similar to the protocol with the QEK algorithm [Subsec. III.1], we test and compare the kernel function (13) for the two cases where the time evolution is either subjected to the global-detuning Hamiltonian $\hat{\mathcal{H}}_{\mathrm{glob}}^{\mathcal{G}}$ (9) or to the local-detuning one $\hat{\mathcal{H}}_{\mathrm{loc}}^{\mathcal{G}}$ (10). We denote the associated kernel functions as $\kappa_{\mathrm{GDQC}}^{\mathrm{glob}}$ and $\kappa_{\mathrm{GDQC}}^{\mathrm{loc}}$, intended for non-attributed and attributed graphs, respectively.
@@
Figure 2: Illustration for the construction of the generalized-distance quantum-correlation feature vector $\bm{\chi}\left[C^{\mathcal{G}}\left(t\right),D^{\mathcal{G}}\right]$ according to the algorithm 1 for the indicated graph. The figure shows for each pair of nodes $(i,j)$ the measured correlation $C^{\mathcal{G}}_{i,j}$ (upper bar) together with the respective distance measure $d_{\mathcal{G}}(i,j)$ (lower bar). Notice how the pairs that are at distance 1 showcase very low values due to the blockade phenomenon. Having chosen $N_{\mathrm{bins}}^{C}=5$ correlation bins, the colors of the upper bars indicate which binning interval they are associated with. For example, there are 16 faint red values within the interval $[0,0.2]$. Within the GDQC framework these values are further discriminated according to the graph distance between the involved nodes. Thus, the 16 faint red values are redistributed according to the four involved graph distance values, leading to a sequence $\{1,7,5,3\}$. All bins combined, this procedure generates the $(N_{\mathrm{bins}}^{C}N_{\mathrm{bins}}^{D})$-dimensional GDQC feature vector $\bm{\chi}$ which, in the present example becomes $\bm{\chi}=(1,3,2,1,0,7,0,0,0,0,5,4,0,0,0,3,2,0,0,0)^{\intercal}/\sqrt{118}$. The presented correlation matrix $C^{\mathcal{G}}_{i,j}=C^{\mathcal{G}}_{i,j}\left(t=1\,\mathrm{\mu s}\right)$ corresponds to the emulation in Fig. 1.

### v2

@@ III Graph Kernel methods for machine learning
In this section we introduce all the kernels used in this work: first, the classical kernels used for benchmarking (III.1); then, the quantum feature kernels, $\mathrm{QEK}$ (III.2) and $\mathrm{GDQC}$ (III.3). Lastly, in Subsec. III.4 we introduce the pooling operations that are employed.
@@
Kernel-based machine learning methods rely on a rigorous mathematical framework grounded in the theory of Reproducing Kernel Hilbert Spaces (RKHS). The foundation is laid by the Moore-Aronszajn Theorem, which guarantees that every positive definite kernel defines a unique RKHS where the kernel acts as the reproducing kernel [4] . Within this space, the Representer Theorem shows that the solution to many regularized risk minimization problems can be written as a finite linear combination of kernel functions evaluated at the training points, thus reducing infinite-dimensional optimization to a tractable finite-dimensional one [43]. Furthermore, Mercer’s Theorem [29] offers a spectral decomposition of continuous, symmetric, positive-definite kernels, interpreting them as inner products in high-dimensional feature spaces [44]. These results together enable the development of powerful nonlinear algorithms, such as Support Vector Machines [15], with solid theoretical guarantees and practical scalability, for various machine learning tasks such as classification and regression. In practical terms these kernel methods allow SVMs to classify non-linearly separable data by implicitly mapping input features into a higher-dimensional space where a linear separator can be found, without the need of an explicit computation of the mapping [44].
III.1 Classical graph kernels
Graph-structured data are challenging for standard kernel methods because graphs do not lie in a vector space with a canonical notion of coordinates or Euclidean geometry. As a result, directly applying generic kernels is often nontrivial, since valid kernels must satisfy symmetry and positive semi-definiteness, and graph similarity is typically defined through discrete, combinatorial comparisons (see Subsec. III.3 for a discussion of properties). This has motivated the development of graph kernels that measure similarity by comparing substructures such as walks, subtrees, and subgraph patterns. By defining an inner product in an implicit feature space that preserves relevant combinatorial information, graph kernels provide a principled and flexible approach to graph machine learning tasks. From the extensive literature on classical graph kernels, we employ the following kernels for benchmarking (practical performance details can be found in App. B):
Random walk kernel
This kernel measures similarity between graphs by counting matching random walks in each graph, thereby capturing detailed structural patterns. Introduced in Ref. [19], it is highly expressive but can be computationally intensive and prone to tottering effects. In our complexity model, the Random Walk kernel scales as $\mathcal{O}(N^{6})$ and is therefore limited to relatively small graphs in practice. In its standard form, the kernel is purely structural and does not natively incorporate node attributes without additional feature engineering.
Graphlet sampling
Graphlet kernels measure similarity between graphs by comparing the frequency of small induced subgraphs, known as graphlets, present in each graph, up to a specified maximum size [38, 48]. In our setting, we enumerate all graphlets up to the chosen size (rather than sampling), which yields exact graphlet frequency vectors and captures local topological features and structural roles of nodes. This improved fidelity comes at higher computational cost and can increase memory consumption as the number of distinct graphlets grows. Exhaustive enumeration of all $k$-node graphlets scales as $\mathcal{O}(N^{k})$. Unless attributes are explicitly included in the graphlet types (for example, by typing graphlets by labels), this baseline primarily captures topology.
Shortest path kernel
Introduced in Ref. [9], the Shortest-Path (SP) kernel is an efficient alternative to random walks. It captures both local and global structure by matching pairs of nodes with identical shortest-path lengths. Its computational complexity scales as $\mathcal{O}(N^{4})$ in our setting. This kernel can incorporate node labels by encoding each node pair as $(\ell(u),\ell(v),D(u,v))$, which enriches the feature space beyond distances alone. However, this also increases the representation dimensionality: the feature size grows from $\mathcal{O}(d_{\mathrm{max}})$ to $\mathcal{O}(|L|^{2}\,d_{\mathrm{max}})$, where $|L|$ denotes the label vocabulary size and $d_{\mathrm{max}}$ the graph diameter.
$k$-core kernel
This kernel measures similarity between graphs by comparing statistics (e.g., distributions or distances) extracted from their $k$-core decomposition. The $k$-core is the maximal induced subgraph in which every node has degree at least $k$, and the sequence of cores as $k$ increases provides a multiscale view of graph connectivity and cohesion. Following Ref. [33], each graph is represented through feature vectors derived from its cores, and graph similarity is computed by applying a base kernel to these structured representations. In our experiments, we use the Shortest-Path kernel as base kernel, thereby combining core-based structural information with distance-aware matching. The $k$-core decomposition can be computed in $\mathcal{O}(\delta_{\mathrm{min}}\,N+M)$ time, where $\delta_{\mathrm{min}}\ll N$ denotes the minimum degeneracy and $M$ the number of edges (the overall cost then includes the cost of the base kernel).
Weisfeiler-Lehman optimal assignment kernel
Building on the Weisfeiler-Lehman (WL) color-refinement procedure, the WL optimal assignment kernel compares two graphs by deriving refined vertex labels across WL iterations and then computing an optimal one-to-one correspondence between their vertex sets under a WL-based base kernel [27]. Concretely, vertices are scored by the number of WL iterations in which their labels match, and the graph-level similarity is obtained by maximizing the total similarity over all bijections between vertices. The resulting assignment kernel can improve upon the WL subtree kernel while remaining valid. Its computational complexity scales as $\mathcal{O}(h\,M+N)$, where $h$ denotes the number of WL iterations. Importantly, the method naturally supports node labels by initializing the refinement with the provided attributes: with labels, nodes start from attribute-derived labels, yielding finer refinements; without labels, all nodes share the same initial label and the kernel captures topology only. Incorporating labels increases the size of the induced label space, which can lead to sparse, high-dimensional representations and weaker generalization on small datasets.
III.2 Quantum evolution kernel ($\mathrm{QEK}$)
@@
First introduced by Henry et al. [24], the quantum evolution kernel algorithm is a quantum-feature kernel implementation [see Eq. (1)] that draws from probability distributions generated from measured observables. Specifically, within the proof-of-principle implementation of Ref. [2], the algorithm draws from the global observable vector $\bm{P}^{\mathcal{G}}=\left(P_{0}^{\mathcal{G}}(t),P_{1}^{\mathcal{G}}(t),P_{2}^{\mathcal{G}}(t),\dots\right)^{\intercal}$ with $P_{k}^{\mathcal{G}}(t)$ being the excitation probability for a fixed number $k$ of excitations at a given time $t$, see Eq. (6). The superscript $\mathcal{G}$ indicates that the time evolution is subjected to a graph-embedded Hamiltonian, cf. Sec. II. To discriminate between two graphs, $\mathcal{G}_{\mu}$ and $\mathcal{G}_{\mu^{\prime}}$, the QEK algorithm implements the following kernel function
$\kappa_{\mathrm{QEK}}(\mathcal{G}_{\mu},\mathcal{G}_{\mu^{\prime}})=\exp\left[-\mu_{0}\,\mathrm{JS}\left(\bm{P}^{\mathcal{G_{\mu}}},\bm{P}^{\mathcal{G_{\mu^{\prime}}}}\right)\right],$ (15)
@@
with the model parameter $\mu_{0}>0$ set to $\mu_{0}=2$ throughout this work. The Jensen-Shannon divergence $\mathrm{JS}\left(\bm{P},\bm{P}^{\prime}\right)$ measures the similarity between the two distributions. By construction, the kernel function (15) ensures positive definiteness.
$\mathrm{QEK}$ post-processing mainly involves turning the measurement results into an excitation probability distribution. With $N_{\mathrm{shots}}$ measurement shots, and a graph of size $N$, forming $\bm{P}^{G}$ takes $\mathcal{O}(N_{\mathrm{shots}}\,N)$ operations (one processes up to $N$ bits per shot). Computing the Jensen-Shannon divergence from this distribution then costs an additional $\mathcal{O}(N)$. Overall, the total post-processing complexity is therefore dominated by $\mathcal{O}(N_{\mathrm{shots}}\,N)$.
@@
Within its proof-of-principle implementation [2], the $\mathrm{QEK}$ kernel (15) has been applied to non-attributed graphs. Correspondingly, the time evolution has been subjected to a global-detuning Hamiltonian $\hat{\mathcal{H}}_{\mathrm{glob}}^{\mathcal{G}}$ (9). When coupled with a support vector machine, the $\mathrm{QEK}$ algorithm has matched the performance of leading state-of-the-art classical graph kernels on several benchmark datasets of small graphs [2]. Thereby, $\mathrm{QEK}$‘s capabilities have been demonstrated both, numerically via emulations as well as on real quantum hardware, on a 32-qubit neutral-atom processor.
@@
In our work, we expand on the previous $\mathrm{QEK}$ implementation [2] twofold. First, and most importantly, we want it to be able to deal with attributed graphs. This is accomplished by changing the Hamiltonian that governs the time evolution from $\hat{\mathcal{H}}_{\mathrm{glob}}^{\mathcal{G}}$ (9) to the local-detuning Hamiltonian $\hat{\mathcal{H}}_{\mathrm{loc}}^{\mathcal{G}}$ (10). As described in Subsec. II.2, the latter Hamiltonian embeds the atomic masses (the node features) into the local detuning $\delta_{i}$. As a result, we obtain an observable vector $\bm{P}^{\mathcal{G}}$ which contains graph information drawn from both, the edges and the node features. Similar to before, the observable vectors are inserted into the kernel function (15). To distinguish the two cases, evolution under $\hat{\mathcal{H}}_{\mathrm{glob}}^{\mathcal{G}}$ and $\hat{\mathcal{H}}_{\mathrm{loc}}^{\mathcal{G}}$, we use the notation $\kappa_{\mathrm{QEK}}^{\mathrm{glob}}$ and $\kappa_{\mathrm{QEK}}^{\mathrm{loc}}$, respectively. Mathematically, switching from $\kappa_{\mathrm{QEK}}^{\mathrm{glob}}$ to $\kappa_{\mathrm{QEK}}^{\mathrm{loc}}$ has profound consequences which can be summarized by proposition 1 which is proven in App. C.
@@
The quantum evolution kernel $\kappa_{\mathrm{QEK}}^{\mathrm{loc}}$ is genuinely more expressive than $\kappa_{\mathrm{QEK}}^{\mathrm{glob}}$ for distinguishing attributed graphs.
@@
The second extension to the $\mathrm{QEK}$ algorithm that we undertake lies in pooling of different time steps. The above implementation draws the excitation probability from a single time step $t$. In Subsec. III.4 we explain how multiple time steps can be combined to accrue information sampled from different regimes of the time evolution.
To gain intuition about the performance of $\mathrm{QEK}$ at later times, we also define a classical baseline. To this end, we exploit the early-time dynamics (see Subsec. II.4) and train the model using only leading-order approximation to the excitation-probability distribution,
$\bm{P}^{\mathcal{G}}_{\mathrm{baseline}}$ $=\left(1-N\Omega^{2}t^{2}/4,N\Omega^{2}t^{2}/4,0,\dots,0\right)^{\intercal},$ (16)
cf. Eqs. (12)-(13). Importantly, the only graph-dependent quantity entering Eq. (16) is the graph size $N=\lvert\mathcal{V}\rvert$. This baseline therefore isolates the contribution of graph size and provides a convenient reference when evaluating $\mathrm{QEK}$ trained on the full quantum data, i.e., outside from the early-time regime. When applying the baseline (16) to the datasets in Sec. IV, we choose $\Omega t=0.25$, which satisfies $N\Omega^{2}t^{2}/4<1$ for all graphs in the datasets.
@@
The excitation probability $P_{k}^{\mathcal{G}}(t)$ (6) used in the $\mathrm{QEK}$ algorithm is a natural observable candidate to capture graph information as it is permutation invariant. Moreover, it is a global observable rendering it a useful tool to study global graph properties. However, if it comes to local graph properties the excitation probability $P_{k}^{\mathcal{G}}(t)$ may not be the most convenient choice. An attempt to generalize the algorithm from the global observables $P_{k}^{\mathcal{G}}(t)$ to observables that capture distributions of local subcomponents of a graph (e.g., nodes or edges), while maintaining permutation invariance, requires the fixing of a canonical ordering for each graph, a costly task known as the graph canonization problem [22]. This potential shortcoming of the $\mathrm{QEK}$ algorithm (15) has motivated the design of the generalized-distance quantum-correlation kernel, see Subsec. III.3.
III.3 Generalized-distance quantum-correlation ($\mathrm{GDQC}$) kernel
@@
In this section, we introduce the generalized-distance quantum-correlation kernel, a new quantum-feature kernel used in this work. Unlike $\mathrm{QEK}$, which is designed to capture global graph information, $\mathrm{GDQC}$ is based on local observables and is therefore better suited to predicting local graph properties.
The $\mathrm{GDQC}$ kernel is inspired by the classical generalized-distance Weisfeiler-Leman (GD-WL) kernel [47]. Following this idea, we combine a graph distance with an additional source of graph information. We denote by $d_{\mathcal{G}}(i,j)$ the distance between nodes $i$ and $j$, and by $D^{\mathcal{G}}$ the corresponding distance matrix with entries $D^{\mathcal{G}}_{i,j}=d_{\mathcal{G}}(i,j)$. In this work, we use the shortest-path distance as the standard metric for characterizing pairwise node distance. This choice offers two key advantages: first, on unweighted graphs it takes integer values, which provides a natural discretization for one of the model’s hyper-parameters and simplifies parameterization. Second, as one of the most well-established graph-theoretic metrics, it benefits from a rich ecosystem of efficient algorithms, making it both principled and computationally practical.
Within the $\mathrm{GDQC}$ framework, the second element is a local observable. Concretely, we use the correlation matrix $C^{\mathcal{G}}(t)$ with entries $C^{\mathcal{G}}_{i,j}(t)=\left\langle\psi_{\mathcal{G}}(t)\big|\hat{n}_{i}\hat{n}_{j}\big|\psi_{\mathcal{G}}(t)\right\rangle$, cf. Eq. (6). More generally, the $\mathrm{GDQC}$ framework can be extended to other local observables, including higher-order correlators.
@@
Algorithm 1 Computation of the $\mathrm{GDQC}$ feature vector $\bm{\chi}\!\left[C^{\mathcal{G}}\!\left(t\right),D^{\mathcal{G}}\right]$ in Eq. (17).
The proposed $\mathrm{GDQC}$ kernel function [cf. Eq. (1)] between two graphs, $\mathcal{G}_{\mu}$ and $\mathcal{G}_{\mu^{\prime}}$, is defined as scalar product,
@@
It involves a graph-dependent $\mathrm{GDQC}$ feature vector $\bm{\chi}\left[C^{\mathcal{G}}\left(t\right),D^{\mathcal{G}}\right]$ whose construction is the core element of the $\mathrm{GDQC}$ framework. Mathematically, it is defined via algorithm 1, and its construction is illustrated in Fig. 2. It relies on a binning procedure where the $N^{2}$ node pairs $\left(i,j\right)$ are binned in two steps: first, according to the value of their correlation-matrix elements $C^{\mathcal{G}}_{i,j}\left(t\right)\in\left[0,1\right]$, and second, according to their distance $d_{\mathcal{G}}(i,j)$. Hereby, the correlation values are binned into $N_{\mathrm{bins}}^{C}\geq 1$ intervals of size $\varDelta_{C}=1/N_{\mathrm{bins}}^{C}$ (bins associated with the upper axis in Fig. 2), while the distances are binned into $N_{\mathrm{bins}}^{D}\geq 1$ intervals of size $\varDelta_{D}=d_{\mathrm{max}}/N_{\mathrm{bins}}^{D}$ where $d_{\mathrm{max}}$ denotes the maximum distance across the entire dataset. Using the integer-valued shortest-path distance (or graph distance), the most natural binning choice is $N_{\mathrm{bins}}^{D}=d_{\mathrm{max}}+1$ bins, i.e. one bin for each possible graph distance (including distance 0) as illustrated on the lower axis in Fig. 2. For the correlation values, we instead treat the binning resolution as a hyperparameter: we analyze the effect of the correlation-bin width $\varDelta_{C}=1/N_{\mathrm{bins}}^{C}$ in Subsec. IV.1. Unless stated otherwise, the results in the main part of Sec. IV use $N_{\mathrm{bins}}^{C}=10$. In combination the two binning procedures create the $(N_{\mathrm{bins}}^{C}N_{\mathrm{bins}}^{D})$-dimensional $\mathrm{GDQC}$ feature vector $\bm{\chi}$ where each element is associated with a specific combination of correlation and distance bins.
Computing the correlation matrix from $N_{\mathrm{shots}}$ measurement outcomes requires $\mathcal{O}(N_{\mathrm{shots}}\,N^{2})$ operations. Computing the all-pairs shortest-path distance matrix with the Floyd-Warshall algorithm [14] costs $\mathcal{O}(N^{3})$. Finally, the binning step is linear in the number of node pairs, and therefore costs $\mathcal{O}(N^{2})$. Overall, the $\mathrm{GDQC}$ post-processing complexity is $\mathcal{O}(N_{\mathrm{shots}}\,N^{2}+N^{3})$. In our experiments, $N_{\mathrm{shots}}>N$, so the dominant term is $\mathcal{O}(N_{\mathrm{shots}}\,N^{2})$.
Following the protocol used for $\mathrm{QEK}$ (Subsec. III.2), we evaluate the kernel in Eq. (17) under two time evolutions: one generated by the global-detuning Hamiltonian $\hat{\mathcal{H}}_{\mathrm{glob}}^{\mathcal{G}}$ (9) and one generated by the local-detuning Hamiltonian $\hat{\mathcal{H}}_{\mathrm{loc}}^{\mathcal{G}}$ (10). We denote the corresponding kernels by $\kappa_{\mathrm{GDQC}}^{\mathrm{glob}}$ and $\kappa_{\mathrm{GDQC}}^{\mathrm{loc}}$. In our experiments, these are used for non-attributed and attributed graphs, respectively.
@@
Figure 2: Illustration for the construction of the generalized-distance quantum-correlation feature vector $\bm{\chi}\left[C^{\mathcal{G}}\left(t\right),D^{\mathcal{G}}\right]$ according to the algorithm 1 for the indicated graph of size $N=7$. The figure shows the measured correlation $C^{\mathcal{G}}_{i,j}$ (upper bar) together with the respective distance measure $d_{\mathcal{G}}(i,j)$ (lower bar) for each of the $N^{2}=49$ node pairs $\left(i,j\right)$. Note the symmetries $C^{\mathcal{G}}_{j,i}=C^{\mathcal{G}}_{i,j}$ and $d_{\mathcal{G}}(j,i)=d_{\mathcal{G}}(i,j)$. Also, notice how the pairs that are at distance 1 showcase very low values due to the blockade phenomenon. Having chosen $N_{\mathrm{bins}}^{C}=5$ correlation bins, the node pairs $\left(i,j\right)$ are binned according to their correlation-matrix values (indicated by the coloring), leading to an intermediate correlation-matrix binning vector $\bm{\chi}^{C}_{\mathrm{bin}}=\left(36,10,2,1,0\right)^{\intercal}$. Next, the node pairs are further binned according to the graph distance $d_{\mathcal{G}}(i,j)$. For example, the 36 elements from the interval $[0,0.2]$ (faint red values) are redistributed according to the four involved graph distance values, leading to a sequence $\{2,14,14,6\}$. All bins combined, this procedure generates the $(N_{\mathrm{bins}}^{C}N_{\mathrm{bins}}^{D})$-dimensional $\mathrm{GDQC}$ feature vector $\bm{\chi}$ which, in the present example becomes $\bm{\chi}=(2,2,2,1,0,14,0,0,0,0,14,4,0,0,0,6,4,0,0,0)^{\intercal}/\sqrt{473}$. The presented correlation matrix $C^{\mathcal{G}}_{i,j}=C^{\mathcal{G}}_{i,j}\left(t=0.4\,\mathrm{\mu s}\right)$ corresponds to the time step indicated in Fig. 1.